OpenStudy (clairebracken1234):

Each pack of hot dogs contains 10 hot dogs, but each pack of hot dog buns contains 8 buns. Phil buys a number of these packs for a barbecue. After the barbecue, Phil finds that he has 4 hot dogs left over. What is the SECOND smallest number of packs of hot dogs he could have bought

jimthompson5910 (jim_thompson5910):

Well if Phil buys 2 backs of hotdogs and 2 packs of buns, then he'll have 20 hotdogs and 16 buns total 20-16 = 4 is the number of leftover hotdogs So 2 packs of each is the smallest number he can buy. But they're looking for the second smallest

jimthompson5910 (jim_thompson5910):

Does it say he has to buy the same number of packs for both?

OpenStudy (clairebracken1234):

No he does not @jim_thompson5910

jimthompson5910 (jim_thompson5910):

ok another combination he could do is 10 packs of hotdogs (10*10 = 100 total) and 12 packs of buns (12*8 = 96 total) 100-96 = 4 left over hotdogs I'm trying other combos to see if there's anything smaller

jimthompson5910 (jim_thompson5910):

6 packs of hotdogs (6*10 = 60) 7 packs of buns (7*8 = 56) 60 - 56 = 4 left over hotdogs

jimthompson5910 (jim_thompson5910):

yeah I think that's the smallest amount he can buy

OpenStudy (clairebracken1234):

6 packs is correct thank u so much ;)) @jim_thompson5910

jimthompson5910 (jim_thompson5910):

no problem